Use the starting balance sheet and statement of cash flows to answer the question. Valley Technology Balance Sheet As of DeceUse the starting balance sheet and statement of cash flows to answer the question. Valley Technology Balance Sheet As of December 31, 2018 (amounts in thousands) Cash 97,000 Accounts Payable 15,000 Accounts Receivable 45,000 Debt 29,000 Inventory 38,000 Other Liabilities 8,000 Property Plant & Equipment, Gross 239,000 Total Liabilities 52,000 Accumulated Depreciation 75,000 Paid-In Capital 73,000 Property Plant & Equipment, Net 164,000 Retained Earnings 236,000 Other Assets 17,000 Total Equity 309,000 Total Assets 361,000 Total Liabilities & Equity 361,000 Valley Technology Statement of Cash Flows January 1 to March 31, 2019 (amounts in thousands) Net Income 8,500 Depreciation 1,500 Decrease (Increase) in Accounts Receivable (100) Decrease (Increase) in Inventory 500 Increase (Decrease) in Accounts Payable (800) Other Adjustments 0 Net Cash Flow from Operating Activities 9,600 Purchase of Property, Plant, & Equipment (5,700) Other Adjustments 0 Net Cash Flow from Investing Activities (5,700) Increase (Decrease) in Debt (1,000) Dividends (200) Other Adjustments 0 Net Cash Flow from Financing Activities (1,200) Net Cash Flow 2,700 What is the value for Total Liabilities & Equity on March 31, 2019? Please specify your answer in the same units as the financial statements.
Answer #1

Total Liabilities & Equity on March 31, 2019:

Accounts payable = 15000 - 800 = 14200

Debt = 29000 - 1000 = 28000

Other Liabilities = 8000

Paid in capital = 73000

Retained Earnings = 236000 + 8500 - 200 = 244300

Total Liabilities & Equity = 14200 + 28000 + 8000 + 73000 + 244300 = 367500